OpenStudy (lieutenantgeneral):

If f(x) = 0.6(3 – x), what is the value of f(–3)?

OpenStudy (***[isuru]***):

u have to plug (-3) instead of "x" in f(x) f(x) = 0.6(3 – x) f( -3) = 0.6 ( 3 - (-3) ) = 0.6 ( 3 + 3) = 0.6 x 6 = 3.6 got it :) ?